# MEDIC: a network for monitoring data quality in collider experiments

### v1

@@ 1 Introduction
Validation of any scientific theory in physics fundamentally depends on the reliability of experimental data. The more complex an experiment is, the harder it becomes to ensure the integrity of that data. Particle detectors represent the pinnacle of engineering and experimental physics, but they also carry the immense challenge of maintaining error free data acquisition. At the scale of experiments such as the Large Hadron Collider (LHC) at CERN (the European Organization for Nuclear Research), where vast amounts of data are collected continuously, a robust and efficient data quality monitoring (DQM) system is indispensable. Each of the experimental apparatus at the LHC [CMS:2008xjf, ATLAS:2008xda, LHCb:2008vvz, ALICE:2008ngc] requires own specific DQM analysis due to their unique and complex multi-component structures [DQM_CMS, DQM_ATLAS_on, DQM_ATLAS_off, DQM_LHCb, DQM_ALICE]. In the following, we will use the CMS detector as a prototypical example whenever required, however all statements made will be general enough to apply to any other experiment.
@@
  • Online monitoring: This is the first step in the DQM chain, where shifters review the live feedback from the detector while the data is being taken. This stage enables rapid identification and resolution of issues. Data rejected at this step will not be processed for further analysis [CMS:2018qpa, ATLAS:2019fst].
  • Offline monitoring: Also referred to as the data certification step, this phase involves a thorough review and certification of the collected data before it is used for physics analysis [Kalsi:20208L].
@@
Given the complexity of the particle detectors, resolving detector-level failures is an inevitable part of any collider operation. Because of the built-in redundancies in the detectors, the relevant physics can often still be extracted even if some local components of the detectors under perform. Nevertheless, as part of good scientific practice, all imperfections in the detector during a run must be recorded. Therefore, an effective DQM system should not only be sensitive to failures but also specific enough to identify the precise source of the issue. The ever increasing complexity of the detectors and the sheer volume of monitoring data necessitate new methods to be implemented in high-energy physics (HEP) monitoring, as the traditional approaches are reaching their limits. In recent years, several machine learning (ML) models have already been introduced to assist the human shifters [pol:hal-03159873, CMSECAL:2023fvz, Brinkerhoff:2025rob, Gavrikov:2025tzc], gradually steering the field of DQM toward increasing automation.
The main objective of this paper is to introduce an end-to-end simulation-driven DQM framework. In a fully developed DQM analysis, readout information from each detector component must ideally be incorporated to capture the full spectrum of possible malfunctions. As an initial step towards that realization, we have modified [^2] At the time of writing this paper this feature has already been included as a part of the standard Delphes installation, but for a detailed outline of the required modification, the reader can follow the Delphes fork at [arghya_delphes]. the Delphes fast simulation framework [deFavereau:2013fsa], to simulate various categories of detector malfunctions. Being a fast simulation, Delphes generates data at the level of reconstructed particle kinematics rather than raw detector signals, the details of which are discussed in the following section. Traditional DQM approaches primarily rely on statistical consistency tests such as the $\chi^{2}$ or the Kolmogorov–Smirnov tests to compare histograms against reference datasets [Piparo:2012gm]. In contrast, ML-driven DQM methods involve supervised classifiers, often convolutional neural networks (CNNs) or autoencoder-based architectures [pol:hal-03159873, CMSECAL:2023fvz], or a hybrid approach of combining both statistical techniques and unsupervised ML [Brinkerhoff:2025rob] or transformer based attention models [Gavrikov:2025tzc] or transfer learning methods [Asres_2025] to identify anomalies in detector read-outs.
@@
One of the major advantages of employing a simulation-driven approach is that the reference run, representing the detector’s ideal, glitch-free condition, can be generated trivially through the same simulation chain. This capability enables a self-consistent and reproducible DQM framework that does not depend on pre-certified data from actual detector runs. Without such a ground-up simulation strategy, one could alternatively inject random errors into certified data and train an ML model to detect anomalies. While these might teach a model to recognize that an anomaly exists, identifying its physical origin would still require intervention from a human expert. The present approach eliminates this need for manual intervention by simulating realistic detector glitches directly, rather than introducing random perturbations. Moreover, this approach is particularly valuable in the context of the High-Luminosity LHC (HL-LHC) upgrade [Apollinari:2015wtw], where substantial detector modifications and new failure modes are expected. Instead of waiting for certified data to retrain DQM models, simulation-based reference datasets can be updated immediately to reflect hardware or configuration changes. This flexibility not only reduces latency in the data certification process (both online and offline) but also enables continuous DQM development, allowing the monitoring algorithms to evolve in parallel with detector upgrades. The complete MEDIC software pipeline, including simulation configurations and anomaly-generation parameter cards for Delphes simulations are available publicly in the GitHub repository [medic]. This open framework allows reproducibility and easy adaptation to other LHC experiments or future colliders.
@@ 2 Detector simulation with irregularities
Ideally, a full fledged DQM analysis requires both reconstructed particle kinematics and highly granular information from the underlying detector subsystems. Although such electronic level detail can also be generated using a full simulation through Geant4 based detector simulators [GEANT4:2002zbu], we kept that computationally intensive approach for future purposes. As a first step toward this broader goal, we adopt Delphes [deFavereau:2013fsa], a fast-simulation framework that provides realistic but higher-level detector responses. While Delphes does not model low-level electronics or detailed shower development, it allows the user to define the geometry, segmentation, and resolution of each detector subsystem such as the tracker, electromagnetic calorimeter (ECAL), and hadronic calorimeter (HCAL) to emulate the behavior of a realistic experimental setup. The simulation chain takes care of effects of the magnetic field and each of the sub-detector resolutions as well as detector smearing to produce a ROOT library for visualization. In this study, the response of the Compact Muon Solenoid (CMS) detector was modeled using the SimpleCalorimeter module implemented in Delphes. We further extended this module by introducing the option to define InsensitiveEtaPhiBins [arghya_delphes], enabling the simulation of localized detector malfunctions by deactivating any specific regions of choice in the calorimeter. This feature allows controlled emulation of potential readout failures, which in real detectors can arise from any unplanned electrical or hardware issues.
@@
Figure 1: Tower eta distribution for different scenarios.
For the present simulation, the ECAL was configured with a granularity of $\Delta\eta\times\Delta\phi=0.02\times 0.02$ in both the barrel ($|\eta|<1.5$) and endcap regions ($1.5<|\eta|<3.0$), while the forward region ($3.0<|\eta|<5.0$) followed the CMS calorimeter segmentation setup [Penzo:2009zz]. The corresponding energy resolution and performance parameters were chosen to be consistent with the official CMS detector parameterizations [^3] The complete Delphes configuration cards, including the ECAL and HCAL resolution parameters, are available in the associated GitHub repository [medic]. [CMS:2013lxn, CMS:2015xaf]. The HCAL was modeled with non-uniform segmentation corresponding to three tower configurations: $5^{\circ}$ towers covering the barrel region ($|\eta|\lesssim 1.6$), $10^{\circ}$ towers for the forward regions ($1.6\lesssim|\eta|\lesssim 4.3$), and $20^{\circ}$ towers for the very forward calorimeter ($4.3<|\eta|\leq 5.0$). The minimum energy thresholds for HCAL towers were set to $1.0$ GeV with SmearTowerCenter set to true to simulate finite spatial resolution effects. By defining InsensitiveEtaPhibBins, one can then switch off any part of choice or even the entirety of the detector components. To illustrate the procedure, $Z\rightarrow ee$ events444with the dataset from http://cp3.irmp.ucl.ac.be/downloads/z_ee.hep.gz. are simulated as an example with both the ECAL and HCAL for the barrel regions of the CMS detector being turned off. For a DQM analysis, one can then use fig.˜1(a) as the reference run and fig.˜1(b) as the anomalous run where no particles are located near the barrel region of the CMS detector.
@@
To generate training and validation data for the present work, we first simulated proton–proton collisions at a center-of-mass energy ($\sqrt{s}$) of $13$ TeV using MadGraph5_aMC@NLO [Alwall:2014hca], with Pythia8 [Bierlich:2022pfr] for parton showering and hadronization. The output particle list was then passed to Delphes for detector-level simulation under the following four configurations:
@@
These controlled modifications, as shown in fig.˜2, simulate distinct classes of detector malfunctions and together form the labeled dataset used to train and validate the MEDIC framework. Note that all four simulations use the same set of particles generated through the MadGraph5_aMC@NLO and Pythia8 pipeline. This is there to ensure that the underlying physics process is identical across scenarios, so MEDIC learns only the glitches in the detector and not differences in event generation process. Although our modification in Delphes [arghya_delphes] allows each individual detector element inside the ECAL and HCAL to be switched off, for this first proof of concept as well as for computational ease, we chose to deactivate sections of the detector as a single unit. The main inspiration of this approach is to demonstrate that the detector glitches can be realistically simulated and that such simulations can and should be incorporated into the development of ML based DQM tools.
@@ 3.1 Architecture
Both the track and tower information sequences are inherently unordered sets of reconstructed objects, with no meaningful physical ordering even in the detector coordinates or in the $30$ track and tower information that are randomly selected. Therefore, to ensure permutation invariance, each collection is first projected into an embedding space via a linear layer and then processed by a transformer encoder based on multi-head self-attention [vaswani2023attentionneed]. While self-attention is permutation-covariant (its output changes consistently under any reordering of the inputs), the permutation invariance is achieved by the subsequent attention-pooling layer, which assigns learnable weights to the embedded objects and aggregates them into a fixed length representation. This pooling step produces a vector of dimension $\left[\mathcal{W},128\right]$, where $128$ is simply a chosen hyperparameter for the embedding dimension. This design guarantees that the final representations are insensitive to any reshuffling of tracks or tower inputs, a crucial property for collider data where object ordering carries no physical meaning. Unlike track and tower features, MET details form a single event-level vector of dimension $3$ with energy, $\eta$ and $\phi$ information. Consequently, MET is processed through a non-linear projection into the same embedding dimension of $128$ as the track and tower encoders.
@@
This architecture, shown in figure 3 therefore integrates permutation-invariant set encoders for tracks and towers, a dedicated embedding pathway for MET, and a convolutional module for classification. This construction yields a model well suited for DQM tasks that can rely on heterogeneous, structured detector information.
@@ 3.2 Cost function & Accuracy metrics
Since the target of MEDIC is to learn a probability distribution, given the $\mathcal{W}$ sets of event data, a Kullback-Leibler (KL) divergence [Joyce2011] between the predicted and target probability distribution is used as the cost function for training. Given the predicted log-probabilities $\log p_{\theta}(y\,|\,x)$ and the target distribution q(y|x)q(y\,|\,x), where $x$ and $y$ denotes the input and output values and $\theta$ denotes the MEDIC hyper-parameters, the per–sample KL divergence used as cost function is defined as
@@
where the sum runs over all $4$ classes of probabilities. In the present implementation of MEDIC [medic], the PyTorch [10.1145/3620665.3640366] function F.kl_div(log_probs, target_probs) is used where it expects the first argument to be log-probabilities and the second to be the target probabilities.
@@
where $c_{i}$ is the true class index and $c$ is the output of the network for a given data $x_{i}$ with $N$ being the total number of data points. While intuitive, this formulation of accuracy does not reflect the full capability of MEDIC network in assigning probabilities for each of the classes. To address this, a modified version of Brier score [^7] Confer [Brier1950] for the original formulation. is also computed, which works as a scoring rule that measures the mean squared error between the predicted probability vector $p_{\theta}(c\,|\,x_{i})$ and the target probability vector $y_{i}$ as
@@
   $L_{\mathrm{val}}=$ accuracy.
@@ 3.4 Training
With the above components in place, the training procedure used for the MEDIC network can be summarized as follow. In this work, a 55–fold cross-validation (or $k=5$) strategy is employed, resulting in an ensemble of five individually trained MEDIC models, each with its own set of parameters $\left\{\theta_{j}\right\}_{j=1}^{5}$ while the training is optimized stochastically using Adam [Kingma2014AdamAM]. For every fold, batch size is fixed to be $64$ and a learning rate of $10^{-3}$ is chosen. To prevent overfitting, the training accuracy defined in (3.2) is continuously monitored, and early stopping is triggered once improvement saturates. The overall training workflow is provided in Algorithm 1.
@@ 4 Results
$30$ $\mathbf{0.897}$ $\mathbf{0.963}$ $\mathbf{0.001}$ $\mathbf{0.903}$ $\mathbf{0.961}$
$40$ $0.891$ $0.962$ $\mathbf{0.001}$ $0.901$ $0.957$
@@
Keeping the network architecture fixed, and employing the same early stopping method as outlined in Algorithm 1, the results for all tested values of $\mathcal{W}$ are summarized in table˜1. From table˜1 it can be observed that MEDIC attains its optimum overall performance at $\mathcal{W}=30$ among the tests that were performed. This behavior can be attributed to the fact that, for a fixed number of hidden layers (and nodes therein): a window that is too small restricts the amount of information available to the model, whereas a window that is too large introduces redundancy and increases the difficulty of learning stable decision boundaries [Hornik1991]. Consequently, for all subsequent discussions including full training curves and confusion matrices, we focus exclusively on the choice $\mathcal{W}=30$, which provides the best empirical performance as observed.
All the experiments were performed on a Linux server featuring dual AMD EPYC $7313$ processors ($64$ threads) and eight NVIDIA RTX A5000 GPUs ($24$ GB GDDR6 memory). All training used CUDA-accelerated PyTorch. For $\mathcal{W}=30$, it took approximately $16$ hours to complete the training with early stopping enabled. The evaluation of metrices during the training process is shown in figure 4.
@@
The confusion matrices and ROC-AUC curves are summarized in figure 5. For the multi-class case, hard predictions as defined in (3.2) is implemented for the confusion matrix and a one-vs-rest approach for ROC-AUC, taking $3000$ samples per class from the test set to get a balanced dataset. For the binary case (normal vs. anomalous), $3000$ normal and $3000$ anomalous samples are used. Although MEDIC is not explicitly tuned for binary classification, it still achieves strong performance with most normal and anomalous events correctly identified. The small number of misclassification reflects the collapsed probability mass from the three anomalous classes into one as mentioned before, yet the network still effectively separates normal from anomalous data.

### v2

# arXiv:2511.18172v2[hep-ex] 27 Feb 2026

@@ 1 Introduction
Validation of any scientific theory in physics fundamentally depends on the reliability of experimental data. The more complex an experiment is, the harder it becomes to ensure the integrity of that data. Particle detectors represent the pinnacle of engineering and experimental physics, but they also carry the immense challenge of maintaining error free data acquisition. At the scale of experiments such as the Large Hadron Collider (LHC) at CERN (the European Organization for Nuclear Research), where vast amounts of data are collected continuously, a robust and efficient data quality monitoring (DQM) system is indispensable. Each of the experimental apparatus at the LHC [19, 1, 8, 3] requires own specific DQM analysis due to their unique and complex multi-component structures [23, 21, 5, 6, 33]. In the following, we will use the CMS detector as a prototypical example whenever required, however all statements made will be general enough to apply to any other experiment.
@@
  • Online monitoring: This is the first step in the DQM chain, where shifters review the live feedback from the detector while the data is being taken. This stage enables rapid identification and resolution of issues. Data rejected at this step will not be processed for further analysis [10, 2].
  • Offline monitoring: Also referred to as the data certification step, this phase involves a thorough review and certification of the collected data before it is used for physics analysis [26].
@@
Given the complexity of the particle detectors, resolving detector-level failures is an inevitable part of any collider operation. Because of the built-in redundancies in the detectors, the relevant physics can often still be extracted even if some local components of the detectors under perform. Nevertheless, as part of good scientific practice, all imperfections in the detector during a run must be recorded. Therefore, an effective DQM system should not only be sensitive to failures but also specific enough to identify the precise source of the issue. The ever increasing complexity of the detectors and the sheer volume of monitoring data necessitate new methods to be implemented in high-energy physics (HEP) monitoring, as the traditional approaches are reaching their limits. In recent years, several machine learning (ML) models have already been introduced to assist the human shifters [31, 4, 18, 24], gradually steering the field of DQM toward increasing automation.
The main objective of this paper is to introduce an end-to-end simulation-driven DQM framework. In a fully developed DQM analysis, readout information from each detector component must ideally be incorporated to capture the full spectrum of possible malfunctions. As an initial step towards that realization, we have modified [^2] At the time of writing this paper this feature has already been included as a part of the standard Delphes installation, but for a detailed outline of the required modification, the reader can follow the Delphes fork at [14]. the Delphes fast simulation framework [22], to simulate various categories of detector malfunctions. Being a fast simulation, Delphes generates data at the level of reconstructed particle kinematics rather than raw detector signals, the details of which are discussed in the following section. Traditional DQM approaches primarily rely on statistical consistency tests such as the $\chi^{2}$ or the Kolmogorov–Smirnov tests to compare histograms against reference datasets [30]. In contrast, ML-driven DQM methods involve supervised classifiers, often convolutional neural networks (CNNs) or autoencoder-based architectures [31, 4], or a hybrid approach of combining both statistical techniques and unsupervised ML [18] or transformer based attention models [24] or transfer learning methods [15] to identify anomalies in detector read-outs.
@@
One of the major advantages of employing a simulation-driven approach is that the reference run, representing the detector’s ideal, glitch-free condition, can be generated trivially through the same simulation chain. This capability enables a self-consistent and reproducible DQM framework that does not depend on pre-certified data from actual detector runs. Without such a ground-up simulation strategy, one could alternatively inject random errors into certified data and train an ML model to detect anomalies. While these might teach a model to recognize that an anomaly exists, identifying its physical origin would still require intervention from a human expert. The present approach eliminates this need for manual intervention by simulating realistic detector glitches directly, rather than introducing random perturbations. Moreover, this approach is particularly valuable in the context of the High-Luminosity LHC (HL-LHC) upgrade [12], where substantial detector modifications and new failure modes are expected. Instead of waiting for certified data to retrain DQM models, simulation-based reference datasets can be updated immediately to reflect hardware or configuration changes. This flexibility not only reduces latency in the data certification process (both online and offline) but also enables continuous DQM development, allowing the monitoring algorithms to evolve in parallel with detector upgrades. The complete MEDIC software pipeline, including simulation configurations and anomaly-generation parameter cards for Delphes simulations are available publicly in the GitHub repository [13]. This open framework allows reproducibility and easy adaptation to other LHC experiments or future colliders.
@@ 2 Detector simulation with irregularities
Ideally, a full fledged DQM analysis requires both reconstructed particle kinematics and highly granular information from the underlying detector subsystems. Although such electronic level detail can also be generated using a full simulation through Geant4 based detector simulators [7], we kept that computationally intensive approach for future purposes. As a first step toward this broader goal, we adopt Delphes [22], a fast-simulation framework that provides realistic but higher-level detector responses. While Delphes does not model low-level electronics or detailed shower development, it allows the user to define the geometry, segmentation, and resolution of each detector subsystem such as the tracker, electromagnetic calorimeter (ECAL), and hadronic calorimeter (HCAL) to emulate the behavior of a realistic experimental setup. The simulation chain takes care of effects of the magnetic field and each of the sub-detector resolutions as well as detector smearing to produce a ROOT library for visualization. In this study, the response of the Compact Muon Solenoid (CMS) detector was modeled using the SimpleCalorimeter module implemented in Delphes. We further extended this module by introducing the option to define InsensitiveEtaPhiBins [14], enabling the simulation of localized detector malfunctions by deactivating any specific regions of choice in the calorimeter. This feature allows controlled emulation of potential readout failures, which in real detectors can arise from any unplanned electrical or hardware issues.
@@
Figure 1: Pedagogic illustration of tower eta distribution for different scenarios.
For the present simulation, the ECAL was configured with a granularity of $\Delta\eta\times\Delta\phi=0.02\times 0.02$ in both the barrel ($|\eta|<1.5$) and endcap regions ($1.5<|\eta|<3.0$), while the forward region ($3.0<|\eta|<5.0$) followed the CMS calorimeter segmentation setup [29]. The corresponding energy resolution and performance parameters were chosen to be consistent with the official CMS detector parameterizations [^3] The complete Delphes configuration cards, including the ECAL and HCAL resolution parameters, are available in the associated GitHub repository [13]. [20, 27]. The HCAL was modeled with non-uniform segmentation corresponding to three tower configurations: $5^{\circ}$ towers covering the barrel region ($|\eta|\lesssim 1.6$), $10^{\circ}$ towers for the forward regions ($1.6\lesssim|\eta|\lesssim 4.3$), and $20^{\circ}$ towers for the very forward calorimeter ($4.3<|\eta|\leq 5.0$). The minimum energy thresholds for HCAL towers were set to $1.0$ GeV with SmearTowerCenter set to true to simulate finite spatial resolution effects. By defining InsensitiveEtaPhibBins, one can then switch off any part of choice or even the entirety of the detector components. To illustrate the procedure, $Z\rightarrow ee$ events444with the dataset from http://cp3.irmp.ucl.ac.be/downloads/z_ee.hep.gz. are simulated purely for pedagogic completeness as an example with both the ECAL and HCAL for the barrel regions of the CMS detector being turned off. This illustrative configuration is not intended to represent the experimental setup used in the present work, but rather to provide a clear visual demonstration of the effect of insensitive detector regions. For a DQM analysis, one can then use fig.˜1(a) as the reference run and fig.˜1(b) as the anomalous run where no particles are located near the barrel region of the CMS detector.
@@
To generate training and validation data for the present work, we first simulated proton–proton collisions at a center-of-mass energy ($\sqrt{s}$) of $13$ TeV using MadGraph5_aMC@NLO [9], with Pythia8 [16] for parton showering and hadronization. The output particle list was then passed to Delphes for detector-level simulation under the following four configurations:
@@
These controlled modifications, as shown in fig.˜2, simulate distinct classes of detector malfunctions and together form the labeled dataset used to train and validate the MEDIC framework. Note that all four simulations, keep the ECAL unaltered and use the same set of particles generated through the MadGraph5_aMC@NLO and Pythia8 pipeline. This is there to ensure that the underlying physics process is identical across scenarios, so MEDIC learns only the glitches in the detector and not differences in event generation process. Although our modification in Delphes [14] allows each individual detector element inside the ECAL and HCAL to be switched off, for this first proof of concept as well as for computational ease, we chose to deactivate sections of the detector as a single unit. The main inspiration of this approach is to demonstrate that the detector glitches can be realistically simulated and that such simulations can and should be incorporated into the development of ML based DQM tools.
@@ 3.1 Architecture
Both the track and tower information sequences are inherently unordered sets of reconstructed objects, with no meaningful physical ordering even in the detector coordinates or in the $30$ track and tower information that are randomly selected. Therefore, to ensure permutation invariance, each collection is first projected into an embedding space via a linear layer and then processed by a transformer encoder based on multi-head self-attention [32]. While self-attention is permutation-covariant (its output changes consistently under any reordering of the inputs), the permutation invariance is achieved by the subsequent attention-pooling layer, which assigns learnable weights to the embedded objects and aggregates them into a fixed length representation. This pooling step produces a vector of dimension $\left[\mathcal{W},128\right]$, where $128$ is simply a chosen hyperparameter for the embedding dimension. This design guarantees that the final representations are insensitive to any reshuffling of tracks or tower inputs, a crucial property for collider data where object ordering carries no physical meaning. Unlike track and tower features, MET details form a single event-level vector of dimension $3$ with energy, $\eta$ and $\phi$ information. Consequently, MET is processed through a non-linear projection into the same embedding dimension of $128$ as the track and tower encoders.
@@
This architecture, shown in figure 3 therefore integrates permutation-invariant set encoders for tracks and towers, a dedicated embedding pathway for MET, and a convolutional module for classification. Therefore, from the standpoint of inference-time computational cost, MEDIC consists of a fixed per-event encoding stage followed by a window-level aggregation and classification stage. For each event within a window of size $\mathcal{W}$, the track and tower branches process a bounded number of objects ($30$ tracks with $7$ features each, $30$ towers with $8$ features each, and a $3$ dimensional MET vector) through linear projections, self-attention, and pooling, yielding a constant-cost event embedding of fixed dimension. These per-event embeddings, together with the MET branch, are stacked into a tensor of shape $\left[128,3,\mathcal{W}\right]$ (cf. the STACK layer in Fig. 3) and passed to the convolutional and pooling layers. As a result, the dominant inference-time cost scales linearly with the window size $\mathcal{W}$ while the per-event processing cost remains constant by construction. This scaling behavior provides a quick estimate of the order-of-magnitude computational resources required when MEDIC is deployed for DQM inference.
@@ 3.2 Cost function & Accuracy metrics
Since the target of MEDIC is to learn a probability distribution, given the $\mathcal{W}$ sets of event data, a Kullback-Leibler (KL) divergence [25] between the predicted and target probability distribution is used as the cost function for training. Given the predicted log-probabilities $\log p_{\theta}(y\,|\,x)$ and the target distribution q(y|x)q(y\,|\,x), where $x$ and $y$ denotes the input and output values and $\theta$ denotes the MEDIC hyper-parameters, the per–sample KL divergence used as cost function is defined as
@@
where the sum runs over all $4$ classes of probabilities. In the present implementation of MEDIC [13], the PyTorch [11] function F.kl_div(log_probs, target_probs) is used where it expects the first argument to be log-probabilities and the second to be the target probabilities.
@@
where $c_{i}$ is the true class index and $c$ is the output of the network for a given data $x_{i}$ with $N$ being the total number of data points. While intuitive, this formulation of accuracy does not reflect the full capability of MEDIC network in assigning probabilities for each of the classes. To address this, a modified version of Brier score [^7] Confer [17] for the original formulation. is also computed, which works as a scoring rule that measures the mean squared error between the predicted probability vector $p_{\theta}(c\,|\,x_{i})$ and the target probability vector $y_{i}$ as
@@
   $L_{\mathrm{val}}=$ KL loss over $\mathcal{D}^{(j)}_{\mathrm{val}}$.
@@ 3.4 Training
With the above components in place, the training procedure used for the MEDIC network can be summarized as follow. In this work, a 55–fold cross-validation (or $k=5$) strategy is employed, resulting in an ensemble of five individually trained MEDIC models, each with its own set of parameters $\left\{\theta_{j}\right\}_{j=1}^{5}$ while the training is optimized stochastically using Adam [28]. For every fold, batch size is fixed to be $64$ and a learning rate of $10^{-3}$ is chosen. To prevent overfitting, the training accuracy defined in (3.2) is continuously monitored, and early stopping is triggered once improvement saturates. The overall training workflow is provided in Algorithm 1.
@@ 4 Results
$30$ $0.897$ $0.963$ $0.001$ $0.903$ $0.961$
$40$ $0.891$ $0.962$ $0.001$ $0.901$ $0.957$
$50$ $0.898$ $0.962$ $0.001$ $0.921$ $0.957$
$60$ $0.906$ $0.964$ $0.0008$ $0.916$ $0.959$
$70$ $0.904$ $0.964$ $0.0006$ $0.916$ $0.956$
$80$ $0.882$ $0.962$ $0.0007$ $0.884$ $0.954$
$90$ $0.895$ $0.962$ $0.0006$ $0.913$ $0.954$
@@
Keeping the network architecture fixed, and employing the same early stopping method as outlined in Algorithm 1, the results for all tested values of $\mathcal{W}$ are summarized in table˜1. A clear improvement is observed when increasing $\mathcal{W}$ from small values up to $\mathcal{W}=30$, indicating that very short windows do not provide sufficient contextual information for reliable classification. For larger window sizes, the performance metrics saturate as accuracy, AUC, and calibration remain stable for $\mathcal{W}>30$ with no systematic improvement or degradation observed up to $\mathcal{W}=90$. We therefore choose $\mathcal{W}=30$ for subsequent analyses as the smallest window size at which near-maximal performance is achieved. This choice balances classification performance with practical considerations, since larger windows increase training and inference cost without providing consistent performance gains. Larger window sizes primarily demonstrate the robustness of the approach rather than defining a distinct optimal regime. Consequently, for all subsequent discussions including full training curves and confusion matrices, we focus exclusively on the choice $\mathcal{W}=30$, to illustrate the performance and robustness of the MEDIC framework.
All the experiments were performed on a Linux server featuring dual AMD EPYC $7313$ processors ($64$ threads) and eight NVIDIA RTX A5000 GPUs ($24$ GB GDDR6 memory). All training used CUDA-accelerated PyTorch. For $\mathcal{W}=30$, it took approximately $16$ hours to complete the training with early stopping enabled. The evaluation of metrices during the training process is shown in figure 4. While the metrics reported in table˜1 quantify training and validation performance, MEDIC is intended to be used purely in inference mode during actual DQM operation. Therefore the relevant computational cost is determined by the per-window inference time and memory footprint. Since events inside a window are processed independently by the track, tower, and MET encoders before aggregation (cf. the STACK layer in fig.˜3), the inference time scales approximately linearly with the window size $\mathcal{W}$, while the memory usage scales linearly with both $\mathcal{W}$ and the fixed number of selected tracks and towers per event. With the input dimensionality capped at a finite number of tracks and towers, which is chosen to be $30$ for the present analysis, the per-event processing cost remains bounded, making the approach compatible with fast or online DQM workflows. This inference-time scaling, rather than the offline training cost, is the relevant consideration for practical deployment.
@@
The confusion matrices and ROC-AUC curves are summarized in figure 5. For the multi-class case, hard predictions as defined in (3.2) is implemented for the confusion matrix and a one-vs-rest approach for ROC-AUC, taking $3000$ samples per class from the test set to get a balanced dataset. For the binary case (normal vs. anomalous), $3000$ normal and $3000$ anomalous samples are used. Although MEDIC is not explicitly tuned for binary classification, it still achieves strong performance with most normal and anomalous events correctly identified. The small number of misclassifications reflects the collapsed probability mass from the three anomalous classes into one as mentioned before, yet the network still effectively separates normal from anomalous data.
@@
To relate the reported performance metrics and accuracy information to practical DQM usage, the network output can be interpreted as an anomaly score evaluated for each sliding window. In an operational setting, alerts would not be triggered by a single anomalous window, but only when $\mathcal{N}$ consecutive windows exceed a chosen anomaly threshold. This requirement helps avoid reacting to short-lived fluctuations while still allowing persistent detector issues to be identified. For example, the binary confusion matrix in fig.˜5(c) indicates a per-window false-positive rate of $\approx 18\%$, which will be reduced exponentially when several consecutive anomalous windows are taken into consideration. In practice, the value of $\mathcal{N}$ would be chosen by the experiment to balance sensitivity against the acceptable false-alert rate, with larger $\mathcal{N}$ reducing accidental alerts at the cost of a longer response time, consistent with a human-in-the-loop DQM procedures. The high AUC and favorable calibration observed for $\mathcal{W}=30$ ensure that such thresholds can be adjusted according to experiment-specific requirements.